September 29, 2021 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. CCAL HDCCAL log book
  4. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
  5. Photon Reconstruction Task Force

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Short term CALWG Items
  2. Long term items
  3. Calorimeter work packages
  4. AI - BCAL Monitoring:
    1. Hydra Real-Time Dashboard
    2. Hydra 24-hr Monitoring Logs

Tentative Agenda

  1. Announcements
  2. Run Operations
    1. FCAL
    2. BCAL
    3. CCAL
  3. Calibration
  4. Monitoring
    1. BCAL LED
      1. Status of Monitoring [2]
    2. BCAL (offline)
    3. FCAL (offline)
      1. pi0 mass shifts when the LED pulser runs at high rate
  5. Reconstruction
  6. Simulations
  7. Action Items
  8. Any other business


Attendees: Mark, Tolga, Varun, JOn, Sean, Malte, Karthik, Elton, Susan, Churamani, Rebecca, Matt, Zisis

  1. Announcements
  2. Run Operations
    1. FCAL: new/refurbished bases installed. PMTs are failing at a higher rate than before (under Colin, there were 3/week) but mostly because of the more aggressive protocols in testing and removing for repairs. The hall is open today so Mark will replace some bases. Malte reported that the group has tested many PMTs and noticed that some when set at 1800 V draw 20 mA (vs nA normally). These will be brought down by 50 or 100 V, which seems to fix the issue. It is thus not advisable to increase the HV to the max recommended, as was being contemplated by PrimEx to balance the gains. This behaviour might be the reason for the scorch marks on that one PMT reported previously. Shift workers: look at the FCAL scalers for signs of problems. Zisis will inform the RC.
    2. BCAL: Elton reported that the pedestals are not as uniform. Sergei could adjust them, but Mark reminded us that these are sensitive to ambient temperature. The decision is to ask Tim to put the screens in front of the crates which may result in more uniform temps, and then evaluate again before adjusting pedestals. When we saw this before, the A/C used to go on every 4 hours, and Facilities people were convinced to change it to every 2 hours. We could also looks at temp sensors (strip charts are hard to analyze).
    3. CCAL
  3. Calibration: no report.
  4. Monitoring
    1. BCAL LED
      1. Varun gave an update on the status of Monitoring by showing the ratio of the LED signals for Up/DN stream SiPMs, when UP/DN LEDs are flashing. The DN ones are not there in the Front Panel (FP) plot. Last time this happened a reboot of the crate was done that seemed to help, but the mechanism is not understood. Mark did a remote reboot this time, but the problem persists. Suspects: hardware or timing.
    2. BCAL (offline): Tolga and Zisis both reported that no data is coming into Hydra. Zisis will ask Kishan and Thomas to increase the monitoring interval from 24 to 48 hours.
    3. FCAL (offline: Susan reported on the pi0 mass shifts when the LED pulser runs at high rate. (Long term item 1). She retraced a conversation in the Sep 20, 2018 CALWG meeting, and blog entry 3595928 and 3597968. She is looking at block number influence. Matt suggested that the mass of the π0 should be looked at as a function of LED intensity and also radius of the FCAL, since the light varies across its face. Mark suggested looking at pedestals (average or event-by-event). This was an initial look and more will be coming.
  5. Reconstruction: Churamani will present an update at our next meeting.
  6. Simulations: no report.
  7. Action Items:
    1. Short: no updates
    2. Long: small updates: 1 has been undertaken by Susan, 3 is being worked on for JEF, 5 has been started by Tegan.
    3. Packages: no updates
  8. Any other business